# Hao Zheng Ma Jiang Chen Tian Chen Chen Xie

### 25 Mar 2026 v1

@@ 4.1 Main Results
We compare against Vevo2 [9], a token-based autoregressive model with disentangled timbre and melody control, where the timbre and melody references share the same clip for singing voice editing, or use separate clips for melody control. Vevo2 is the most direct baseline, as other systems operate under fundamentally different paradigms: in-context learning approaches require manually aligned edit boundaries and are restricted to local segments, while SoulX-Singer relies on precise character-level timestamps that are either impractical to obtain or would grant an unrealistic advantage if sourced from curated datasets.
As shown in Table 2, YingMusic-Singer consistently outperforms Vevo2 across all six modification types under both Melody Control and Sing Edit settings in PER, F0-CORR, and VS, demonstrating strong adherence to both the reference melody and modified lyrics. The intelligibility gap is most pronounced on Trans and Mix tasks, indicating that reconstructing a substantially different phoneme sequence while preserving melody is inherently challenging. Note that PER on Mix tasks may be further inflated by ASR hallucinations on code-switched utterances. Vevo2's incomplete melody disentanglement tends to reduce intelligibility and introduce hallucinations, whereas YingMusic-Singer's unified IPA tokenization and GRPO-based lyric adherence optimization maintain robustness even under these extreme conditions. F0-CORR further distinguishes the two systems: benefiting from CKA alignment and GRPO, YingMusic-Singer maintains consistently high correlation across all tasks and languages, whereas Vevo2 fluctuates considerably, suggesting less robust melody control. For SIM, Vevo2 benefits from its multi-stage architecture, where an autoregressive LLM handles melody and content generation while a dedicated CFM focuses on timbre reconstruction, effectively easing speaker modeling. YingMusic-Singer instead adopts a single-stage CFM that jointly models all factors, prioritizing architectural simplicity for practical deployment. However, Vevo2 achieves higher SIM but simultaneously exhibits degraded PER and F0-CORR. In practice, faithfully rendering modified lyrics while preserving the original melodic structure remains the primary concern in lyric editing. As shown in Table 3, YingMusic-Singer consistently achieves higher N-MOS and M-MOS than Vevo2 across both tasks and languages. The strong subjective scores also indicate that GRPO optimization does not overfit to the reward models but generalizes to human perception. Vevo2 receives lower ratings overall with notably higher variance, and listeners report perceptible artifacts such as unfaithful lyric rendering and melodic misalignment in its outputs, suggesting less robust generation quality.
@@ 5 Conclusion
We present YingMusic-Singer, a melody-controllable singing voice editing model that synthesizes from a timbre reference, a melody-providing singing clip, and modified lyrics without manual alignment. Through curriculum training and GRPO-based reinforcement learning, YingMusic-Singer achieves superior melody preservation and lyric adherence on LyricEditBench, the first comprehensive benchmark we introduce for this task, demonstrating strong potential for practical end-to-end singing voice editing.

### 09 Apr 2026 v2

@@ 4.1 Main Results
We compare against Vevo2 [zhang2025vevo2unifiedcontrollableframework], a token-based autoregressive model with disentangled timbre and melody control, where the timbre and melody references share the same clip for singing voice editing, or use separate clips for melody control. Vevo2 is the most direct baseline, as other systems operate under fundamentally different paradigms: in-context learning approaches require manually aligned edit boundaries and are restricted to local segments, while SoulX-Singer relies on precise character-level timestamps that are either impractical to obtain or would grant an unrealistic advantage if sourced from curated datasets.
As shown in Table 2, YingMusic-Singer-Plus consistently outperforms Vevo2 across all six modification types under both Melody Control and Sing Edit settings in PER, F0-CORR, and VS, demonstrating strong adherence to both the reference melody and modified lyrics. The intelligibility gap is most pronounced on Trans and Mix tasks, indicating that reconstructing a substantially different phoneme sequence while preserving melody is inherently challenging. Note that PER on Mix tasks may be further inflated by ASR hallucinations on code-switched utterances. Vevo2's incomplete melody disentanglement tends to reduce intelligibility and introduce hallucinations, whereas YingMusic-Singer-Plus's unified IPA tokenization and GRPO-based lyric adherence optimization maintain robustness even under these extreme conditions. F0-CORR further distinguishes the two systems: benefiting from CKA alignment and GRPO, YingMusic-Singer-Plus maintains consistently high correlation across all tasks and languages, whereas Vevo2 fluctuates considerably, suggesting less robust melody control. For SIM, Vevo2 benefits from its multi-stage architecture, where an autoregressive LLM handles melody and content generation while a dedicated CFM focuses on timbre reconstruction, effectively easing speaker modeling. YingMusic-Singer-Plus instead adopts a single-stage CFM that jointly models all factors, prioritizing architectural simplicity for practical deployment. However, Vevo2 achieves higher SIM but simultaneously exhibits degraded PER and F0-CORR. In practice, faithfully rendering modified lyrics while preserving the original melodic structure remains the primary concern in lyric editing. As shown in Table 3, YingMusic-Singer-Plus consistently achieves higher N-MOS and M-MOS than Vevo2 across both tasks and languages. The strong subjective scores also indicate that GRPO optimization does not overfit to the reward models but generalizes to human perception. Vevo2 receives lower ratings overall with notably higher variance, and listeners report perceptible artifacts such as unfaithful lyric rendering and melodic misalignment in its outputs, suggesting less robust generation quality.
@@ 5 Conclusion
We present YingMusic-Singer-Plus, a melody-controllable singing voice editing model that synthesizes from a timbre reference, a melody-providing singing clip, and modified lyrics without manual alignment. Through curriculum training and GRPO-based reinforcement learning, YingMusic-Singer-Plus achieves superior melody preservation and lyric adherence on LyricEditBench, the first comprehensive benchmark we introduce for this task, demonstrating strong potential for practical end-to-end singing voice editing.